If secx = cos^2x then the value of sin^4 x - 2sin^3 x + sin^2 x is

Grade:11

1 Answers

Ankit Jaiswal
7 years ago
Hi Akshay,
                    the answer to your question is 0
    sec x=cos2x
\Rightarrow1/cos x=cos2x          (sec x=1/cos x)
\Rightarrow1=cos3x
\Rightarrowcos x=1                     (\sqrt[3]{1} =1 and not = -1)
so x = 0\degree satisfies the codition
and hence 
sin x=sin0\degree=0
and hence 
sin4x – 2sin3x +sin2x
=0-2*0+0
=0
I hope this might help thanks